# Coarse-Grained Boltzmann Generators

### v1

@@ 3.3 The CG-BG Workflow
After training, the learned PMF $U_{\eta}$ defines the target energy for importance sampling, rather than serving as a force field for MD integration. Let $q_{\theta}(\mathbf{R})$ denote the density induced by the trained flow model. Importance weights are computed as
@@
Provided $U_{\eta}$ accurately approximates the true PMF on the support of $q_{\theta}$, this procedure gives unbiased estimates under $p(\mathbf{R})$.
The reliability of importance reweighting is quantified by the effective sample size (ESS) (Kish, 1965),
$\mathrm{ESS}=\frac{\left(\sum_{i=1}^{B}w(\mathbf{R}_{i})\right)^{2}}{\sum_{i=1}^{B}w(\mathbf{R}_{i})^{2}}.$ (18)
The ESS depends critically on the overlap between the generative density $q_{\theta}$ and the target density defined by $U_{\eta}$. In practice, machine learning potentials may exhibit unphysical extrapolation outside the training domain, and generative models may occasionally produce high-energy artifacts, both of which can lead to weight degeneracy. To improve robustness, we apply a weight clipping strategy (Tan et al., 2025a; Gloy and Olsson, 2025; Moqvist et al., 2025) to truncate statistical outliers before computing expectations (See §G.2). The complete CG-BG training and sampling pipeline is summarized in Fig. 1.
@@ 4 Experiments
We evaluate CG-BGs on the Müller–Brown (MB) potential and alanine dipeptide (capped alanine). Additional experimental details including architectures are provided in §C and §D. CG-BG samples are generated and reweighted following the algorithms described in §F.
Datasets. For both systems, we construct unbiased and biased datasets for training and evaluation. Biased datasets are generated using enhanced sampling methods to accelerate exploration of configurational space. For the MB system, datasets are generated via Langevin dynamics (§B.1), with umbrella sampling (Torrie and Valleau, 1977) applied in the biased setting to facilitate transitions between metastable basins. For alanine dipeptide, we generate datasets using both explicit and implicit solvent models (§B.2). Explicit solvent datasets, used for training, are produced using a classical force field (Lindorff-Larsen et al., 2010) and include long unbiased trajectories as well as biased simulations from well-tempered metadynamics (Barducci et al., 2008). Implicit solvent datasets are generated with the same force field combined with a generalized Born solvent model, using two parameterizations (OBC1 and OBC2) (Onufriev et al., 2004). The explicit solvent trajectories are further coarse-grained using either a Heavy Atom mapping (Fig. 3a), which retains all heavy atoms, or a Core Beta mapping (Fig. 4a), which preserves the backbone and $C_{\beta}$ atoms. Full details on dataset generation and enhanced sampling procedures are provided in §B.
Baselines. Unlike previous BG work (Klein and Noé, 2024; Tan et al., 2025a, b), which often treats implicit solvent simulations as ground truth, we use explicit solvent simulations as reference and regard implicit solvent models as baselines. For alanine dipeptide, CG reduces the effective dimensionality to either 10 beads (Heavy Atom) or 6 beads (Core Beta). To quantify the computational gain via coarse-graining, we use the full 22-atom system as a baseline.
Metrics. We report ESS, Jenson-Shannon (JS) divergence, and PMF error. JS divergence is computed between the sampled and reference dihedral angle free energy profiles. The PMF error is defined as the squared distance between the negative logarithms of the sampled and reference densities, which gives more weight to low-density regions compared to JS divergence (Plainer et al., 2025; Durumeric et al., 2024). Energy histograms and free energy profiles of $\phi$ dihedral are shown in the main text, while additional results, including $\psi$ dihedral free energy profiles, Ramachandran plots (Ramachandran, 1963), bond length distributions, and weight clipping ablations, are provided in §G.
@@ 4.1 Recovering Equilibrium Distributions from Biased and Unbiased Data
We first demonstrate that CG-BGs inherit the importance reweighting capability of atomistic BGs, enabling recovery of unbiased equilibrium statistics from flow models trained on either biased or unbiased trajectories.
@@
For the MB system (Fig. 2a–b), coarse-graining corresponds to projection onto the $x$-coordinate, yielding an analytically exact reference $U(x)=-k_{B}T\ln\int\exp\!\big(-\beta u(x,y)\big)\,dy$ (Fig. 2c–d). For alanine dipeptide, we use the Heavy Atom mapping, which removes hydrogens and solvent degrees of freedom (Fig. 3a). Throughout §4.1 and §4.2, the PMF is learned from rapidly converged datasets using ESFM (§3.2).
Figure 2: CG-BGs on the MB potential. (a) Two-dimensional MB potential energy surface (functional form in §B). (b) Marginal probability density along the $x$ coordinate. (c) Free energy profiles before and after reweighting for CG-BGs, where flow is trained on unbiased data, compared with the exact solution and MD reference. (d) Same as (c), but for flow trained on biased data.
@@
As shown in Fig. 2c and Fig. 3c, despite training on long unbiased datasets and using highly expressive models, the raw flow proposals deviate from the MD reference. These discrepancies arise from low-quality samples generated by the flow model, a limitation intrinsic to Boltzmann Emulators that cannot be systematically corrected without reweighting. After reweighting, however, CG-BGs successfully recover equilibrium free energy profiles in close agreement with MD references (Fig. 2c and Fig. 3b-c). The reweighted distributions accurately reproduce relative basin populations and capture transition regions. This is further illustrated by the Ramachandran plots (Fig. 7). Unreweighted proposals exhibit noisy samples across transition regions and metastable basins (Fig. 7b), which are assigned low importance weights and effectively filtered out after reweighting (Fig. 7c). Quantitative metrics in Tab. 2 confirm close agreement between reweighted samples and target equilibrium ensemble.
@@
We further evaluate CG-BGs trained on biased or short non-equilibrium trajectories (umbrella sampling for MB and 10 ns WT-MetaD with $\gamma=1.5$ for alanine dipeptide), reflecting realistic settings where long unbiased simulations are infeasible. As shown in Fig. 2d, Fig. 3d and Fig. 7c, while the flow proposals exhibit larger deviations from the MD reference, importance reweighting consistently recovers accurate equilibrium statistics (Tab. 2).
Figure 3: CG-BGs on alanine dipeptide (Heavy Atom). (a) Heavy Atom mapping. (b) Potential energy distributions under the learned PMF before and after reweighting, compared with the MD reference. (c) $\phi$ dihedral free energy profile before and after reweighting for CG-BGs, where flow is trained on 500 ns unbiased data, alongside the MD reference. (d) Same as (c), but for flow trained on a 10 ns WT-MetaD ($\gamma=1.5$) dataset.
@@
Notably, after reweighting, CG-BGs outperform implicit solvent MD baselines (Tab. 2), highlighting the advantage of learning PMFs from explicit solvent simulations. While implicit solvent models perform reasonably well for alanine dipeptide, prior work has reported drastic discrepancies for more complex systems (Chen et al., 2021). This constitutes an improvement over previous BG approaches, which rely on implicit solvent models for reweighting and are therefore fundamentally limited by solvent approximation error. In other words, such models can at best achieve the accuracy of implicit solvent baselines (Tab. 2). These results show that CG-BGs generate CG equilibrium samples consistent with the target distribution, without requiring long unbiased MD simulations, and offer a practical route to correct systematic biases in existing Boltzmann Emulators.
@@ 4.2 Effect of Coarse-Graining Resolution on Accuracy and Efficiency
We next examine how the choice of CG resolution affects both modeling accuracy and computational efficiency. To this end, we consider a coarser representation of alanine dipeptide, the Core Beta mapping (Fig. 4a).
Despite the reduced resolution, CG-BGs trained with the Core Beta mapping remain capable of recovering equilibrium statistics after reweighting (Fig. 4c–d and Fig. 10), with quantitative metrics reported in Tab. 2. While performance remains strong for alanine dipeptide, coarser mappings are known to induce a gradual loss of accuracy for more complex systems (Görlich and Zavadlav, 2025). This degradation arises from increased degeneracy in the CG representation. For a given CG coordinate $\mathbf{R}$, many atomistic microstates $\mathbf{r}_{i}$} satisfy $\Xi(\mathbf{r}_{i})=\mathbf{R}$ while exerting different projected forces $\mathcal{F}_{\mathrm{proj}}(\mathbf{r}_{i})$. As a result, the conditional mean force $\mathbb{E}_{p(\mathbf{r}\mid\mathbf{R})}\!\left[\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\right]$ exhibits higher variance under coarser mappings, introducing irreducible uncertainty and making the force matching problem intrinsically more challenging.
Figure 4: CG-BGs on alanine dipeptide (Core Beta). (a) Core Beta mapping. (b) Potential energy distributions under the learned PMF before and after reweighting, compared with the MD reference. (c) $\phi$ dihedral free energy profile before and after reweighting for CG-BGs, where flow is trained on 500 ns unbiased data, alongside the MD reference. (d) Same as (c), but for flow trained on a 10 ns WT-MetaD ($\gamma=1.5$) dataset .
@@
We further benchmark the computational cost of CG-BGs across different resolutions (Tab. 1). As expected, the Core Beta mapping substantially reduces both training and inference time compared to the Heavy Atom and full atomistic representations, with particularly large gains at inference due to cheaper Jacobian evaluations. Note that the reported all atom baseline generates only solute coordinates; generating full configurations with explicit solvent, which would be the proper reference, is computationally infeasible. Overall, these results highlight the favorable accuracy–efficiency trade-off of CG-BGs. Additional comparisons with prior work are provided in Tab. 4.
Table 1: Training and inference time across different CG mappings. Inference times correspond to $10^{4}$ generated samples. All Atom denotes generating the full solute configuration without solvent.
Stage Core Beta Heavy Atom All Atom
Training 0.83h 1.39h 3.93h
Inference 0.84min 3.46min 13.92min
Total 0.84h 1.45h 4.16h
@@
Table 2: Quantitative comparison for alanine dipeptide across CG resolutions and baseline models. CG-BGs results are reported after reweighting; Biased denotes flow trained on WT-MetaD datasets, while $\text{PMF}_{U}$ indicates PMFs learned from long unbiased MD data.
@@
Heavy Atom 0.0075 $\pm$ 0.0001 $0.2493\pm 0.0126$ $0.2325\pm 0.0006$
Heavy Atom (Biased) $0.0090\pm 0.0002$ $0.2758\pm 0.0109$ $0.1936\pm 0.0005$
Core Beta $0.0079\pm 0.0001$ $0.2519\pm 0.0136$ 0.2510 $\pm$ 0.0006
Core Beta (Biased) $0.0081\pm 0.0001$ 0.2325 $\pm$ 0.0093 $0.2239\pm 0.0005$
Heavy Atom ($\text{PMF}_{U}$) $0.0218\pm 0.0002$ $0.3611\pm 0.0138$ $0.2222\pm 0.0006$
Implicit Solvent Baselines (Generalized Born Models)
GB (OBC1) $0.0157\pm 0.0002$ $0.3709\pm 0.0092$
GB (OBC2) $0.0182\pm 0.0002$ $0.4028\pm 0.0095$
4.3 Simulation-Free Evaluation of Learned PMFs
Beyond equilibrium sampling, CG-BGs also provide a practical tool for evaluating learned PMFs without running MD. The flow model generates CG configurations from a proposal distribution approximating the equilibrium ensemble, and importance reweighting maps these samples to the Boltzmann distribution induced by a given PMF. As a result, equilibrium observables for any candidate PMF can be estimated using a single set of generated samples and associated weights, without additional simulations. In contrast to traditional validation pipelines that require long MD runs for each model, CG-BGs enable rapid, one-shot benchmarking of CG potentials.
We leverage this capability to benchmark PMFs trained under different data regimes, comparing models learned from long unbiased MD trajectories ($\text{PMF}_{U}$) with those trained on a rapidly converged biased dataset ($\text{PMF}_{B}$). Quantitative metrics (Tab. 2) computed from reweighted samples provide immediate feedback on model accuracy, while visual diagnostics including probability densities(Fig. 5a), dihedral free energy profiles (Fig. 5b), and Ramachandran plots (Fig. 9) qualitative validation against atomistic references. Consistent with prior observations (Chen et al., 2026; Görlich and Zavadlav, 2025), $\text{PMF}_{U}$ fails to capture the correct metastable population of the $\phi$ angle, whereas $\text{PMF}_{B}$ shows improved agreement.
While demonstrated here for CG models, this simulation-free evaluation paradigm naturally extends to atomistic MLPs. CG-BGs therefore serve not only as efficient equilibrium sampling methods, but also as diagnostic tools for rapid comparison and validation of learned energy models.
Figure 5: Simulation-free benchmarking of learned CG PMFs using CG-BGs (Heavy Atom). (a) Probability density of the $\phi$ dihedral angle after reweighting with PMFs trained on unbiased ($\text{PMF}_{U}$) and rapidly converged biased datasets ($\text{PMF}_{B}$), compared with the MD reference and flow proposal (trained on unbiased data). (b) Corresponding $\phi$ dihedral free energy profiles after reweighting, alongside the MD reference and flow proposal.

### v2

@@ 3.3 The CG-BG Workflow
After training, the learned PMF $U_{\eta}$ defines the target energy for importance sampling, rather than as a force field for MD integration. Let $q_{\theta}(\mathbf{R})$ denote the density induced by the trained flow model. Importance weights are computed as
@@
Provided $U_{\eta}$ accurately approximates the true PMF on the support of $q_{\theta}$, reweighting gives unbiased estimates under $p(\mathbf{R})$ in the idealized setting where the PMF is exact. In practice, the learned PMF introduces approximation bias.
The reliability of importance reweighting is quantified by the normalized effective sample size (ESS) (Kish, 1965),
$\mathrm{ESS}=\frac{1}{B}\frac{\left(\sum_{i=1}^{B}w(\mathbf{R}_{i})\right)^{2}}{\sum_{i=1}^{B}w(\mathbf{R}_{i})^{2}}.$ (18)
where $B$ denotes the number of generated samples. The normalized ESS takes values in $(0,1]$, with larger values indicating better overlap between the generative density $q_{\theta}$ and the target density defined by $U_{\eta}$. In practice, machine learning potentials may exhibit unphysical extrapolation outside the training domain, and generative models may occasionally produce high-energy artifacts, both of which can lead to weight degeneracy. To improve robustness, we apply a weight clipping strategy (Tan et al., 2025a; Gloy and Olsson, 2025; Moqvist et al., 2025) to truncate statistical outliers before computing expectations (See §G.4). The complete CG-BG training and sampling pipeline is summarized in Fig. 1.
Figure 2: CG-BGs on alanine dipeptide. (a–e) Heavy Atom mapping results. (a) Heavy Atom mapping. (b) Potential energy distribution under the learned PMF for flow trained on 500 ns unbiased data, before and after reweighting, compared with MD reference. (c) $\phi$ dihedral free energy profile for the same model. (d) Energy distribution for flow trained on 10 ns WT-MetaD ($\gamma=1.5$) data. (e) Corresponding $\phi$ dihedral free energy profile. (f–j) Core Beta mapping results are shown in the second row with the same structure as (a–e).
Figure 3: Ramachandran plots of alanine dipeptide. (a) MD reference. (b,c) Heavy Atom mapping: reweighted Ramachandran distributions. In (b), the flow model is trained on 500 ns unbiased data; in (c), it is trained on 10 ns WT-MetaD ($\gamma=1.5$) data. (d,e) Core Beta mapping: same training setups as (b,c). Unreweighted Ramachandran plots are provided in Fig. 8.
@@ 4 Experiments
We evaluate CG-BGs on the Müller–Brown (MB) potential and three alanine peptide systems, including alanine dipeptide (Ac-Ala-NHMe, 22 atoms), alanine tripeptide (Ac-Ala3-NHMe, 42 atoms), and alanine hexapeptide (Ac-Ala6-NHMe, 72 atoms). Additional experimental details including architectures are provided in §C and §D. CG-BG samples are generated and reweighted following the algorithms described in §F.
Datasets. For all systems, we construct unbiased and biased datasets for training and evaluation. Biased datasets are generated using enhanced sampling methods to accelerate exploration. For the MB system, data are generated via Langevin dynamics (§B.1), with umbrella sampling (Torrie and Valleau, 1977) used in the biased setting to improve transitions between metastable basins. For peptide systems, datasets are generated using both explicit and implicit solvent models (§B.2). Explicit solvent data are produced using a classical force field (Lindorff-Larsen et al., 2010) and include long unbiased MD trajectories as well as biased simulations obtained via well-tempered metadynamics (Barducci et al., 2008). Implicit solvent data are generated using the same force field in combination with a generalized Born model under different parameterizations (OBC1 and OBC2) (Onufriev et al., 2004). For explicit solvent simulations, configurations are further coarse-grained using either a Heavy Atom mapping (Fig. 2a and Fig. 4a), which retains all heavy atoms, or a Core Beta mapping (Fig. 2f and Fig. 4f), which retains backbone atoms and the $C_{\beta}$ position. Full details on dataset generation and enhanced sampling procedures are provided in §B.
Baselines. Unlike previous BG work (Klein and Noé, 2024; Tan et al., 2025a, b), which often treats implicit solvent simulations as reference, we use explicit solvent simulations as the primary reference and treat empirical implicit solvent models as baselines. We additionally report results from atomistic BGs, including TarFlow and ECNF++ (Tan et al., 2025b) trained on implicit solvent simulation data.
Metrics. We report ESS, Jensen-Shannon (JS) divergence, and PMF error. JS divergence is computed between the sampled and reference dihedral angle free energy profiles. The PMF error is defined as the squared distance between the negative logarithms of the sampled and reference densities, placing additional emphasis on low-probability regions compared to JS divergence (Plainer et al., 2025; Durumeric et al., 2024). Energy histograms and free energy profiles of $\phi$ dihedral are shown in the main text, while additional results, including $\psi$ dihedral free energy profiles, Ramachandran plots (Ramachandran, 1963), bond length distributions, and weight clipping ablations, are provided in §G.
@@ 4.1 Recovering Equilibrium Distributions from Biased and Unbiased Data
We first demonstrate that CG-BGs inherit the importance reweighting capability of atomistic BGs, enabling recovery of equilibrium statistics from flow models trained on either biased or unbiased trajectories.
@@
For the MB system (Fig. 6), coarse-graining corresponds to projection onto the $x$-coordinate, yielding an analytically exact reference $U(x)=-k_{B}T\ln\int\exp\!\big(-\beta u(x,y)\big)\,dy$ (Fig. 6c–d). For peptides, we use two CG mappings depending on system size: the Heavy Atom mapping for alanine dipeptide and tripeptide (Fig. 2a and Fig. 4a), and the Core Beta mapping for hexapeptide (Fig. 4f). Throughout §4.1 and §4.2, the PMF is learned from rapidly converged datasets using ESFM (§3.2).
@@
As shown in Fig. 2c, Fig. 4c, 4h and Fig. 6c, despite training on long unbiased datasets and using expressive models, the raw flow proposals deviate from the MD reference. These discrepancies arise from low-quality samples generated by the flow model, a limitation intrinsic to Boltzmann Emulators that cannot be systematically corrected without reweighting. After reweighting, CG-BGs successfully recover equilibrium free energy profiles in close agreement with MD references. The reweighted distributions accurately reproduce relative basin populations and match the reference distribution in transition regions. This is further illustrated by the Ramachandran plots (Fig. 3, Fig. 4e and Fig. 4j). Unreweighted proposals exhibit noisy samples across transition areas and metastable basins (Fig. 8 and Fig. 9), which are assigned low importance weights and effectively filtered out after reweighting. Quantitative metrics in Tab. 2 confirm close agreement between reweighted samples and target equilibrium ensemble.
@@
We further evaluate CG-BGs trained on biased or short non-equilibrium trajectories (umbrella sampling for MB and 10 ns WT-MetaD with $\gamma=1.5$ for alanine dipeptide), reflecting realistic settings where long unbiased simulations are infeasible. As shown in Fig. 2 and Fig. 6f, while the flow proposals exhibit larger deviations from the MD reference, importance reweighting consistently recovers accurate equilibrium statistics (Tab. 2).
Figure 4: CG-BGs on alanine tripeptide and hexapeptide. (a–e) Alanine tripeptide (Heavy Atom mapping): (a) CG representation, (b) potential energy distribution, (c) $\phi$ free energy profile, (d) MD reference Ramachandran plot, and (e) reweighted Ramachandran distribution. (f–j) Alanine hexapeptide (Core Beta mapping): (f) CG representation, (g) potential energy distribution, (h) $\phi$ free energy profile, (i) MD reference Ramachandran plot, and (j) reweighted Ramachandran distribution. Unreweighted Ramachandran plots are shown in Fig. 9. The reported free energy profiles and Ramachandran plots correspond to the dihedral pair biased during WT-MetaD: the second pair (of three) for alanine tripeptide and the third pair (of six) for alanine hexapeptide, counting from the N-methyl terminus.
@@
Notably, after reweighting, CG-BGs outperform implicit solvent MD baselines (Tab. 2), highlighting the advantage of learning PMFs from explicit solvent simulations. While implicit solvent models perform reasonably well for alanine dipeptide, we observe increasing differences for tripeptide (Fig. 4c) and hexapeptide (Fig. 4h), consistent with previous observations for more complex molecular systems (Chen et al., 2021). This constitutes an improvement over atomistic BG approaches, which rely on implicit solvent models for reweighting and are therefore fundamentally limited by solvent approximation error. In other words, atomistic BGs can at best achieve the accuracy of implicit solvent baselines (Tab. 2 and Tab. 7). These results show that CG-BGs generate CG equilibrium samples consistent with the target distribution, without requiring long unbiased MD simulations, and offer a practical route to correct systematic biases in existing Boltzmann Emulators.
Table 1: Training and inference time across different CG mappings on alanine dipeptide. Inference times correspond to $10^{4}$ generated samples. All Atom denotes generating the full solute configuration without solvent.
Stage Core Beta Heavy Atom All Atom
Training 0.45h 0.80h 2.55h
Inference 0.95min 3.78min 14.91min
Total 0.47h 0.86h 2.80h
@@ 4.2 Effect of Coarse-Graining Resolution on Accuracy and Efficiency
We next examine how the choice of CG resolution affects both sampling quality and computational efficiency. To this end, we consider a coarser Core Beta mapping for alanine dipeptide (Fig. 2f) and tripeptide (Fig. 7a). Despite the reduced resolution, CG-BGs trained with the Core Beta mapping remain capable of recovering equilibrium statistics after reweighting (Fig. 2h, 2j and Fig. 7c), with quantitative metrics reported in Tab. 2 and Tab. 3.
Compared to the Heavy Atom mapping, the lower-dimensional Core Beta representation generally yields higher ESS, indicating improved overlap between the flow proposal and the target distribution. This is expected, as generative modeling and importance sampling become easier in lower-dimensional spaces. Nevertheless, after reweighting, the resulting equilibrium statistics are generally less accurate than those obtained with the Heavy Atom mapping. A likely explanation is the increased degeneracy introduced by coarse-graining. For a given CG coordinate $\mathbf{R}$, many atomistic microstates $\mathbf{r}_{i}$} satisfy $\Xi(\mathbf{r}_{i})=\mathbf{R}$ while exerting different projected forces $\mathcal{F}_{\mathrm{proj}}(\mathbf{r}_{i})$. As a result, the conditional mean force $\mathbb{E}_{p(\mathbf{r}\mid\mathbf{R})}\!\left[\mathcal{F}_{\mathrm{proj}}(\mathbf{r})\right]$ exhibits larger variance under coarser mappings, increasing the difficulty of accurately learning the PMF through force matching. This effect is expected to become more pronounced as molecular complexity increases (Görlich and Zavadlav, 2025).
@@
We further benchmark the computational cost of CG-BGs across different resolutions (Tab. 1). As expected, the Core Beta mapping substantially reduces both training and inference time compared to the Heavy Atom and full atomistic representations, with particularly large gains at inference due to cheaper Jacobian evaluations. Note that the reported all atom baseline generates only solute coordinates; generating full configurations with explicit solvent, which would be the proper reference, is computationally infeasible. Overall, the results illustrate a trade-off between statistical efficiency and representation fidelity: coarser mappings improve proposal quality and computational efficiency, while finer mappings show more accurate equilibrium estimates after reweighting. Additional comparisons with prior work are provided in Tab. 5.
4.3 Simulation-Free Evaluation of Learned PMFs
Beyond equilibrium sampling, CG-BGs provide a form of amortized equilibrium benchmarking for learned PMFs. Whereas conventional BGs use a learned proposal distribution to estimate observables under a fixed target energy, the same proposal distribution can also be reused to evaluate and compare multiple candidate PMFs through importance reweighting. Once a sufficiently accurate proposal has been learned, equilibrium observables under different PMFs can be estimated from a single set of generated samples, without performing additional simulations.
Specifically, the flow model generates CG configurations from a proposal distribution approximating the equilibrium ensemble, and importance reweighting maps these samples to the Boltzmann distribution induced by a given PMF. This enables rapid, simulation-free assessment of candidate CG potentials, in contrast to traditional validation pipelines that require separate MD simulations for each model.
We leverage this capability to compare PMFs trained under different data regimes, contrasting models learned from long unbiased MD trajectories ($\mathrm{PMF}_{U}$) with those trained on a rapidly converged biased dataset ($\mathrm{PMF}_{B}$). Quantitative metrics from reweighted samples (Tab. 2) provide a direct measure of model accuracy, while dihedral free energy profiles and Ramachandran plots enable visual comparison with atomistic references. Consistent with previous observations (Chen et al., 2026; Görlich and Zavadlav, 2025), $\mathrm{PMF}_{U}$ (Fig. 5) fails to recover the correct metastable populations along $\phi$, whereas $\mathrm{PMF}_{B}$ exhibits improved agreement.
Although demonstrated here for CG models, the same principle naturally extends to atomistic machine learning potentials. More generally, generated configurations can be reused to compare multiple learned energy functions without additional simulations, making CG-BGs a practical tool for rapid model validation.
@@ 4.3 Simulation-Free Evaluation of Learned PMFs
Table 2: Quantitative comparison for alanine dipeptide across CG resolutions and baseline models. CG-BGs results are reported after reweighting; Biased denotes flow trained on WT-MetaD datasets, while $\text{PMF}_{U}$ indicates PMFs learned from long unbiased MD data. Flow proposal results are provided in Tab. 7.
@@
Heavy Atom 0.0048(1) 0.2005(63) 0.5112(4)
Heavy Atom (Biased) 0.0063(1) 0.2277(66) 0.4115(4)
Core Beta 0.0052(1) 0.2210(65) 0.5528(4)
Core Beta (Biased) 0.0057(1) 0.2093(58) 0.4818(4)
Heavy Atom ($\text{PMF}_{U}$) 0.0050(1) 0.2131(68) 0.5196(4)
Implicit Solvent Baselines
GB (OBC1) 0.0157(2) 0.3709(92)
GB (OBC2) 0.0182(2) 0.4028(95)
Figure 5: Simulation-free benchmarking of learned CG PMFs using CG-BGs on alanine dipeptide (Heavy Atom). (a) $\phi$ free energy profile after reweighting with PMFs trained on unbiased ($\text{PMF}_{U}$) and rapidly converged biased datasets ($\text{PMF}_{B}$), compared with the MD reference and flow proposal (trained on unbiased data). (b) Ramachandran plot reweighted using $\mathrm{PMF}_{U}$.
Table 3: Quantitative comparison for alanine tripeptide and hexapeptide across CG resolutions and baseline models. CG-BGs results are reported after reweighting. Flow proposal results are provided in Tab. 8 and Tab. 9.
Model Alanine Tripeptide Alanine Hexapeptide
JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$) JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$)
Core Beta 0.0060(1) 0.2112(51) 0.4212(5) 0.0100(1) 0.3646(81) 0.1231(3)
Heavy Atom 0.0056(1) 0.1957(52) 0.3201(4)
Implicit Solvent Baselines
GB (OBC2) 0.0932(3) 1.0274(65) 0.1652(3) 1.8401(70)